# Unified Kraft Break at $\sim\,$6500 K: A Newly Identified Single-Star Obliquity Transition Matches the Classical Rotation Break

### v1

@@ Abstract
facilities: WIYN/NEID, PFS/Magellansoftware: numpy (Oliphant, 2006; Walt et al., 2011; Harris et al., 2020), matplotlib (Hunter, 2007), pandas (McKinney, 2010), scipy (Virtanen et al., 2020), MESA (Paxton et al., 2011, 2013, 2015, 2018, 2019)
1 Introduction
@@
The mass of a star largely determines how it generates energy, and how this energy is transported. Relatively low core temperatures in low-mass cool dwarfs ($0.5-1.3\,\,M_{\sun}$) favor hydrogen burning through the proton-proton chain, with energy carried outward through a deep convective envelope. In hotter, more massive dwarfs, the CNO cycle dominates hydrogen burning in a convective core, while the overlying envelope is predominantly radiative. These structural differences have been supported by multiple observational tracers. For example, cool dwarfs show envelope-driven mixing that depletes surface lithium (Boesgaard & Tripicco, 1986; Hobbs & Pilachowski, 1986; Deliyannis & Pinsonneault, 1997; Sestito & Randich, 2005). Where asteroseismology is observationally available, this convection stochastically excites solar-like oscillations (Kjeldsen & Bedding, 1995; Chaplin & Miglio, 2013; Silva Aguirre et al., 2013). Hotter F stars, with shallower convective envelopes, exhibit far shorter convective turnover and damping timescales in their pulsations. Stars which are hotter still possess negligible convective envelopes, and display only classical pulsations (e.g. $\delta$ Scuti/roAp) driven by the $\kappa$ mechanism (Cox, 1980).
@@
A much more easily accessible observational manifestation of this structural shift is the sharp drop in rotation rates at mid-F spectral types: the well-known Kraft break [^1] Explicit evidence for a sharp drop in rotation near F5 long predates Kraft’s work (e.g., Struve & Elvey 1931; Westgate 1934). (Kraft, 1967). A sufficiently thick convective envelope in cool stars sustains a global magnetic dynamo, whose efficiency is described by the dimensionless Rossby number, operationally defined as the ratio between the rotational period and the convective turnover timescale (Noyes et al., 1984; Brun & Browning, 2017). Low Rossby numbers in cool stars are generally associated with higher magnetic activity, manifesting as enhanced spots, frequent flares, and strong chromospheric/coronal emission (e.g. Saar & Brandenburg, 1999; Pizzolato et al., 2003; Mamajek & Hillenbrand, 2008; Reiners et al., 2009; Vidotto et al., 2014; Stelzer et al., 2016; Newton et al., 2017; Feinstein et al., 2020; See et al., 2023; Mathur et al., 2025). Magnetized stellar winds extract angular momentum and slow the rotation of stars as they age (e.g. Weber & Davis, 1967; Skumanich, 1972; Kawaler, 1989; Barnes, 2003; van Saders & Pinsonneault, 2013). By contrast, with thinner convective envelopes, magnetic braking is much less efficient in hot stars, which remain rapidly rotating on the main sequence.
@@
Independently, exoplanet observations have revealed a transition in the stellar spin-orbit angle, the angle between the stellar rotation axis and the planetary orbital axis (hereafter obliquity), among host stars: hot Jupiters orbiting hotter stars more likely exhibit large misalignments, whereas those around cooler stars are usually well aligned (e.g., Schlaufman, 2010; Winn et al., 2010a; Albrecht et al., 2012, 2021; Knudstrup et al., 2024). This is often attributed to efficient tidal obliquity damping (Albrecht et al., 2012), via dissipation of tidally excited inertial waves within convective envelopes (Ogilvie & Lin, 2007; Lai, 2012; Li & Winn, 2016; Saunders et al., 2024), and, where relevant, by resonances between the dynamical tides and stellar oscillations (Zanazzi et al., 2024; Zanazzi & Chiang, 2025), whereas such damping is inefficient in hot, radiative-envelope stars.
@@
Because both phenomena reflect a structural break separating cool and hot stars, it has long been assumed that the rotation and obliquity transitions coincide. Whether they do has remained unclear. Historically, both the rotation (e.g., Kraft, 1967; Noyes et al., 1984; van Saders & Pinsonneault, 2013; Avallone et al., 2022) and obliquity (e.g., Winn et al., 2010a; Albrecht et al., 2012; Winn & Fabrycky, 2015; Triaud, 2018; Albrecht et al., 2022; Knudstrup et al., 2024) literatures have been content to adopt notional thresholds without fitting them statistically, leading to somewhat disparate breaks: a mid-F (i.e. $M_{\star}=1.2$-$1.3\,\,M_{\sun}$) divider, or the equivalent in color or temperature (such as BV=0.3B\!-\!V=0.3, $G_{\rm BP}\!-\!G_{\rm RP}=0.55$-0.60, or $T_{\rm eff}\sim 6500\ \mathrm{K}$: Iben 1967; do Nascimento et al. 2000; van Saders et al. 2019) for rotation, and $T_{\rm eff}=6000$-$6300$ K (often 6100 or 6250 K, as reviewed by Albrecht et al. 2022) for obliquity, without a formal turning-point inference or quoted uncertainties.
@@
Recent work has sharpened one side of this picture. Using a carefully curated sample of nearby single F-dwarfs, Beyer & White (2024) found a sharp rotation transition centered at $T_{\rm eff}=6550$ K (width of 200 K), arguing that the commonly cited $6100$ K obliquity transition is $450$ K cooler than this value. Although the obliquity sample (dominated by hot-Jupiter hosts) and the rotation sample (composed of F stars within 33 pc) do differ in $\left[{\rm Fe}/{\rm H}\right]$ (medians $0.04^{+0.20}_{-0.16}$ vs. $0.15^{+0.19}_{-0.18}$ ) — which can affect the Kraft-break location by altering envelope opacity (and thus convective-zone depth and magnetic-braking efficiency) — a $\sim\!0.2$ dex metalicity offset is expected to shift the break only by less than $100$ K (Spalding & Winn, 2022): far smaller than the observed $450$ K gap.
@@
In this work, we supply one: we present a consistent statistical determination of the transition points in both stellar $v\sin{i_{\star}}$ (rotation) and $\lambda$ (obliquity), enabling a rigorous comparison of like with like. We apply a uniform “turning-point” fitting methodology to an updated, single-star rotation sample based on Beyer & White (2024), and confirm a sharp rotation transition at $T_{\rm eff}={6510^{+97}_{-127}}$ K. Applying the same procedure to a hot-Jupiter Rossiter-McLaughlin (RM, Rossiter 1924; McLaughlin 1924) sample, we recover an obliquity transition consistent with commonly quoted values, $T_{\rm eff}={6105^{+123}_{-133}}$ K. Thus, our baseline analysis reproduces a similar but slightly smaller $405$ K offset to that noted by Beyer & White (2024). Crucially, when we remove binaries from the obliquity sample so that both probes refer to single stars, the inferred obliquity break shifts upward to $T_{\rm eff}={6447^{+85}_{-119}}$ K, which is statistically indistinguishable from the rotation break ($6510^{+97}_{-127}$} K), differing by only $0.41\,\sigma$. That is to say, a single mid-F boundary at $\sim 6500$ K consistently describes both the rotation (Kraft) break and the spin-orbit alignment transition, once binarity is controlled for.
2 Sample construction
2.1 Stellar Rotation, $v\sin{i_{\star}}$
@@
To statistically identify the turning point of the Kraft rotation break, we assembled a sample of 405 systems with $T_{\rm eff}$ and $v\sin{i_{\star}}$ measurements (panels c and d of Figure 1), following the general selection and vetting criteria described in Section 2 of Beyer & White (2024), but with minor modifications aimed at expanding the sample and minimizing the influence of binary/multiple-star systems.
@@
We compiled all bright F-type stars within 33.33 pc of the Sun that have reliable parallax measurements, using the ADQL query described in Section 2 of Beyer & White (2024). The primary source of $T_{\rm eff}$  is teff_gspphot; if unavailable, we adopt teff_gsspec or teff_esphs [^2] https://gaia.aip.de/metadata/gaiadr3/astrophysical_parameters/. If none of these are available, $T_{\rm eff}$  values from the Geneva–Copenhagen Survey (Nordström et al., 2004) and the stellar property catalog of rapidly rotating Sun-like stars by Schröder et al. (2009) were supplemented. As for $v\sin{i_{\star}}$  the Gaia DR3 line-broadening measurements (vbroad) were adopted as a proxy (Frémat et al., 2023). When vbroad was unavailable, we instead used $v\sin{i_{\star}}$ values from Nordström et al. (2004) and Schröder et al. (2009). Following Beyer & White (2024), we removed young stars (with ages less than 200 Myr) and evolved stars (with $\log g_{\star}<3.75$), as these may exhibit anomalous rotation rates. We identified systems with stellar companions via multiple steps. Firstly, binaries and binary candidates were identified by cross-matching our targets with the Geneva–Copenhagen Survey (Nordström et al., 2004), where entries marked with an asterisk in the $f_{d}$ (confirmed or suspected binaries) or $f_{b}$ (spectroscopic binaries) flags were considered binary or potential binary systems. Next, we identified stellar companions following the methods described in El-Badry et al. (2021), but without the $G<18$ limitation in order to include faint stellar companions. Finally, stars with Renormalized Unit Weight Error (RUWE, Lindegren et al. 2018, 2021) >> 1.4 were classified as potential binary or multiple-star systems. The final sample comprises 405 stars with $T_{\rm eff}$ and $v\sin{i_{\star}}$ measurements, including 182 mature, single main-sequence stars, 206 binary or multiple-star systems (see panels c and d of Figure 1), and 17 young or evolved stars, which are excluded from our analysis.
@@
Figure 1: The $\lambda$ and $v\sin{i_{\star}}$  distributions, along with derived $T_{\rm eff}$  boundaries. Panel a: $\lambda$ distribution for hot-Jupiter systems with confirmed or candidate stellar companions. The $T_{\rm eff}$  boundary of $6105^{+123}_{-133}$} K is marked with a black dashed line, with shaded 1 $\sigma$ interval. Panel b: Same as the top panel, but for single-star hot-Jupiter systems with $T_{\rm eff}$  boundary = $6447^{+85}_{-119}$} K. Panel c: $v\sin{i_{\star}}$  distribution for single stars, compiled following the updated prescription described in Beyer & White (2024). The resulting $T_{\rm eff}$  boundary is $6510^{+97}_{-127}$} K, consistent with that derived from the single-star $\lambda$ distribution. Panel d: same as panel c, but for binary and multiple-star systems.
2.2 Stellar Obliquity, $\lambda$
@@
To determine the boundary of the obliquity transition, we focus on systems that would be affected by tidal realignment, if operative. Concretely, we select hot Jupiters with mass ratios $3\times 10^{-4}\leq m_{\rm p}/M_{\star}\leq 2\times 10^{-3}$ (Rusznak et al., 2025) on close-in orbits with $a/R_{\star}<10$, and with measured stellar obliquities. Stellar masses, planetary masses, and $a/R_{\star}$ are taken from the Planetary Systems Composite Parameters table (PSCompPars [^3] https://exoplanetarchive.ipac.caltech.edu/cgi-bin/TblView/nph-tblView?app=ExoTbls&config=PSCompPars, NASA Exoplanet Archive 2019) at NASA Exoplanet Archive (Christiansen et al., 2025); when a value is missing there, we adopt the entry from the Encyclopaedia of exoplanetary systems [^4] https://exoplanet.eu/catalog/ as a fallback.
@@
Among stellar-obliquity measurement techniques, the Rossiter-McLaughlin method is the most widely used and provides the majority of $\lambda$ measurements (Winn & Fabrycky, 2015; Triaud, 2018; Albrecht et al., 2022). Because different techniques introduce different selection effects (e.g., spot-crossing is more sensitive to aligned systems, while gravity-darkening is usually significant only for highly misaligned systems, Siegel et al. 2023; Albrecht et al. 2022), we retain only RM-based measurements, giving priority in the following order: classical RM, Doppler shadow (Albrecht et al., 2007; Collier Cameron et al., 2010; Zhou et al., 2016; Johnson et al., 2017), reloaded RM (Cegla et al., 2016), and RM Revolution (Bourrier et al., 2021). For all retained systems, we adopt the sky-projected obliquity $\lambda$ and $T_{\rm eff}$  from TEPCat [^5] https://www.astro.keele.ac.uk/jkt/tepcat/obliquity.html,
@@
Stellar multiplicity has long been theorized as a potential driver of spin-orbit misalignments (Holman et al., 1997; Wu & Murray, 2003; Fabrycky & Tremaine, 2007; Naoz et al., 2012; Batygin, 2012). Recent obliquity studies confirm that warm Jupiters tend to be aligned in single-star systems (Rice et al., 2022; Wang et al., 2024; Espinoza-Retamal et al., 2023) yet are frequently (and sometimes famously, e.g., HD 80606; Winn et al. 2009; TIC 241249530; Gupta et al. 2024) misaligned in binaries; massive planets show a similar single-versus-binary contrast (Rusznak et al., 2025). Motivated by this, we identify and flag binaries or higher-order multiples in our obliquity sample using Schwarz et al. (2016) and Fontanive & Bardalez Gagliuffi (2021). Similar to subsection 2.1, we examined the binarity of each system following the method of El-Badry et al. (2021), but extended it to include faint nearby stars ($G>18$). We also classified systems with RUWE >> 1.4 as binary candidates. Moreover, we only include systems with stellar ages >> 200 Myr and $\log g_{\star}$ >> 3.75 to exclude young and evolved stars, which is consistent with the construction of the $v\sin{i_{\star}}$  sample.
@@
Figure 3: The $\lambda$ distributions for hot-Jupiter ($3\times 10^{-4}\leq m_{\rm p}/M_{\star}\leq 2\times 10^{-3}$, $a/R_{\star}$<<10), sub-Saturn ($3\times 10^{-5}\leq m_{\rm p}/M_{\star}\leq 3\times 10^{-3}$), warm-Jupiter ($3\times 10^{-4}\leq m_{\rm p}/M_{\star}\leq 2\times 10^{-3}$, $a/R_{\star}$>>11), compact multiplanet (period ratio of adjacent planets, $P_{i+1}/P_{i}<6$), massive-planet systems ($2\times 10^{-3}\leq m_{\rm p}/M_{\star}$), and eccentric Jupiters ($3\times 10^{-4}\leq m_{\rm p}/M_{\star}\leq 2\times 10^{-3}$, $e>0.1$), along with the Kraft break and its associated uncertainty derived in this work. In terms of $T_{\rm eff}$  coverage, only the hot-Jupiter systems have sufficient measurements (>10>10) in the hot-star regime. Sub-Saturns with nearby companions are marked in blue. The predicted relation between $T_{\rm eff}$  and $\lambda$ from secular resonance crossing (Petrovich et al., 2020; Dugan et al., 2025) is shown as the light-blue shaded region.
3 $T_{\rm eff}$  Boundary Determination

### v2

@@ Abstract
facilities: WIYN/NEID, PFS/Magellansoftware: numpy (oliphant2006guide; walt2011numpy; harris2020array), matplotlib (hunter2007matplotlib), pandas (mckinney2010data), scipy (virtanen2020scipy), MESA (mesa_paper_1; mesa_paper_2; mesa_paper_3; mesa_paper_4; mesa_paper_5)
I Introduction
@@
The mass of a star largely determines how it generates energy, and how this energy is transported. Relatively low core temperatures in low-mass cool dwarfs ($0.5-1.3\,\,M_{\sun}$) favor hydrogen burning through the proton-proton chain, with energy carried outward through a deep convective envelope. In hotter, more massive dwarfs, the CNO cycle dominates hydrogen burning in a convective core, while the overlying envelope is predominantly radiative. These structural differences have been supported by multiple observational tracers. For example, cool dwarfs show envelope-driven mixing that depletes surface lithium (Boesgaard1986; Hobbs1986; Deliyannis1997; Sestito2005). Where asteroseismology is observationally available, this convection stochastically excites solar-like oscillations (Kjeldsen1995; Chaplin2013; Silva2013). Hotter F stars, with shallower convective envelopes, exhibit far shorter convective turnover and damping timescales in their pulsations. Stars which are hotter still possess negligible convective envelopes, and display only classical pulsations (e.g. $\delta$ Scuti/roAp) driven by the $\kappa$ mechanism (Cox1980).
@@
A much more easily accessible observational manifestation of this structural shift is the sharp drop in rotation rates at mid-F spectral types: the well-known Kraft break [^1] Explicit evidence for a sharp drop in rotation near F5 long predates Kraft’s work (e.g., Struve1931; Westgate1934). (kraft1967break). A sufficiently thick convective envelope in cool stars sustains a global magnetic dynamo, whose efficiency is described by the dimensionless Rossby number, operationally defined as the ratio between the rotational period and the convective turnover timescale (noyes_rotation_1984; brun_magnetism_2017). Low Rossby numbers in cool stars are generally associated with higher magnetic activity, manifesting as enhanced spots, frequent flares, and strong chromospheric/coronal emission (e.g. saar_time_1999; pizzolato_stellar_2003; mamajek_improved_2008; reiners_evidence_2009; vidotto_stellar_2014; stelzer_path_2016; newton_halpha_2017; Feinstein2020; see_further_2023; mathur_magnetic_2025). Magnetized stellar winds extract angular momentum and slow the rotation of stars as they age (e.g. weber_AM_1967; skumanich_timescales_1972; kawaler_rotational_1989; barnes_gyrochronology_2003; vansaders_fast_2013). By contrast, with thinner convective envelopes, magnetic braking is much less efficient in hot stars, which remain rapidly rotating on the main sequence.
@@
Independently, exoplanet observations have revealed a transition in the stellar spin-orbit angle, the angle between the stellar rotation axis and the planetary orbital axis (hereafter obliquity), among host stars: hot Jupiters orbiting hotter stars more likely exhibit large misalignments, whereas those around cooler stars are usually well aligned (e.g., Schlaufman2010; Winn2010; Albrecht2012; Albrecht2021; Knudstrup2024). This is often attributed to efficient tidal obliquity damping (Albrecht2012), via dissipation of tidally excited inertial waves within convective envelopes (OgilvieANDLin2007; Lai2012; Li2016; saunders_efficient_2024), and, where relevant, by resonances between the dynamical tides and stellar oscillations (zanazzi2024damping; Zanazzi2025), whereas such damping is inefficient in hot, radiative-envelope stars.
@@
Because both phenomena reflect a structural break separating cool and hot stars, it has long been assumed that the rotation and obliquity transitions coincide. Whether they do has remained unclear. Historically, both the rotation (e.g., kraft1967break; noyes_rotation_1984; vansaders_fast_2013; Avallone2022) and obliquity (e.g., Winn2010; Albrecht2012; WinnFabrycky2015; Triaud2018; Albrecht2022; Knudstrup2024) literature have been content to adopt notional thresholds, treating their locations as well established, despite rarely determining them through formal statistical fitting and, in many cases, without identifying the boundary explicitly in terms of stellar effective temperature. As a result, the quoted transition locations vary across studies, leading to somewhat disparate breaks: a mid-F (i.e. $M_{\star}=1.2$-$1.3\,\,M_{\sun}$) divider, or the equivalent in color or temperature (such as BV=0.3B\!-\!V=0.3, $G_{\rm BP}\!-\!G_{\rm RP}=0.55$-0.60, or $T_{\rm eff}\sim 6500\ \mathrm{K}$: Iben1967Comparison_of_Theory_with_Observation; donascinmento_lithium_2000; vansaders_forward_2019) for rotation, and $T_{\rm eff}=6000$-$6300$ K (often 6100 or 6250 K, as reviewed by Albrecht2022) for obliquity, without a formal turning-point inference or quoted uncertainties.
@@
Recent work has sharpened one side of this picture. Using a carefully curated sample of nearby single F-dwarfs, Beyer2024 found a sharp rotation transition centered at $T_{\rm eff}=6550$ K (width of 200 K), arguing that the commonly cited $6100$ K obliquity transition is $450$ K cooler than this value. Although the obliquity sample (dominated by hot-Jupiter hosts) and the rotation sample (composed of F stars within 33 pc) do differ in $\left[{\rm Fe}/{\rm H}\right]$ (medians $0.04^{+0.20}_{-0.16}$ vs. $0.15^{+0.19}_{-0.18}$ ) — which can affect the Kraft-break location by altering envelope opacity (and thus convective-zone depth and magnetic-braking efficiency) — a $\sim\!0.2$ dex metalicity offset is expected to shift the break only by less than $100$ K (Spalding2022): far smaller than the observed $450$ K gap.
@@
In this work, we supply one: we present a consistent statistical determination of the transition points in both stellar $v\sin{i_{\star}}$ (rotation) and $\lambda$ (obliquity), enabling a rigorous comparison of like with like. We apply a uniform “turning-point” fitting methodology to an updated, single-star rotation sample based on Beyer2024, and confirm a sharp rotation transition at $T_{\rm eff}={6510^{+97}_{-127}}$ K. Applying the same procedure to a hot-Jupiter Rossiter-McLaughlin (RM, Rossiter1924; McLaughlin1924) sample, we recover an obliquity transition consistent with commonly quoted values, $T_{\rm eff}={6105^{+123}_{-133}}$ K. Thus, our baseline analysis reproduces a similar but slightly smaller $405$ K offset to that noted by Beyer2024. Crucially, when we remove binaries from the obliquity sample so that both probes refer to single stars, the inferred obliquity break shifts upward to $T_{\rm eff}={6447^{+85}_{-119}}$ K, which is statistically indistinguishable from the rotation break ($6510^{+97}_{-127}$} K), differing by only $0.41\,\sigma$. That is to say, a single mid-F boundary at $\sim 6500$ K consistently describes both the rotation (Kraft) break and the spin-orbit alignment transition, once binarity is controlled for.
II Sample construction
II.1 Stellar Rotation, $v\sin{i_{\star}}$
@@
To statistically identify the turning point of the Kraft rotation break, we assembled a sample of 405 systems with $T_{\rm eff}$ and $v\sin{i_{\star}}$ measurements (panels c and d of Figure 1), following the general selection and vetting criteria described in Section 2 of Beyer2024, but with minor modifications aimed at expanding the sample and minimizing the influence of binary/multiple-star systems.
@@
We compiled all bright F-type stars within 33.33 pc of the Sun that have reliable parallax measurements, using the ADQL query described in Section 2 of Beyer2024. The primary source of $T_{\rm eff}$  is teff_gspphot; if unavailable, we adopt teff_gsspec or teff_esphs [^2] https://gaia.aip.de/metadata/gaiadr3/astrophysical_parameters/. If none of these are available, $T_{\rm eff}$  values from the Geneva–Copenhagen Survey (Nordstrom2004) and the stellar property catalog of rapidly rotating Sun-like stars by Schroder2009 were supplemented. As for $v\sin{i_{\star}}$  the Gaia DR3 line-broadening measurements (vbroad) were adopted as a proxy (Fremat2023). When vbroad was unavailable, we instead used $v\sin{i_{\star}}$ values from Nordstrom2004 and Schroder2009. Following Beyer2024, we removed young stars (with ages less than 200 Myr) and evolved stars (with $\log g_{\star}<3.75$), as these may exhibit anomalous rotation rates. We identified systems with stellar companions via multiple steps. Firstly, binaries and binary candidates were identified by cross-matching our targets with the Geneva–Copenhagen Survey (Nordstrom2004), where entries marked with an asterisk in the $f_{d}$ (confirmed or suspected binaries) or $f_{b}$ (spectroscopic binaries) flags were considered binary or potential binary systems. Next, we identified stellar companions following the methods described in Badry2021, but without the $G<18$ limitation in order to include faint stellar companions. Finally, stars with Renormalized Unit Weight Error (RUWE, Lindegren2018; Lindegren2021) >> 1.4 were classified as potential binary or multiple-star systems. The final sample comprises 405 stars with $T_{\rm eff}$ and $v\sin{i_{\star}}$ measurements, including 182 mature, single main-sequence stars, 206 binary or multiple-star systems (see panels c and d of Figure 1), and 17 young or evolved stars, which are excluded from our analysis.
@@
Figure 1: The $\lambda$ and $v\sin{i_{\star}}$  distributions, along with derived $T_{\rm eff}$  boundaries. Panel a: $\lambda$ distribution for hot-Jupiter systems with confirmed or candidate stellar companions. The $T_{\rm eff}$  boundary of $6105^{+123}_{-133}$} K is marked with a black dashed line, with shaded 1 $\sigma$ interval. Panel b: Same as the top panel, but for single-star hot-Jupiter systems with $T_{\rm eff}$  boundary = $6447^{+85}_{-119}$} K. Panel c: $v\sin{i_{\star}}$  distribution for single stars, compiled following the updated prescription described in Beyer2024. The resulting $T_{\rm eff}$  boundary is $6510^{+97}_{-127}$} K, consistent with that derived from the single-star $\lambda$ distribution. Panel d: same as panel c, but for binary and multiple-star systems.
II.2 Stellar Obliquity, $\lambda$
@@
To determine the boundary of the obliquity transition, we focus on systems that would be affected by tidal realignment, if operative. Concretely, we select hot Jupiters with mass ratios $3\times 10^{-4}\leq m_{\rm p}/M_{\star}\leq 2\times 10^{-3}$ (Rusznak2025) on close-in orbits with $a/R_{\star}<10$, and with measured stellar obliquities. Stellar masses, planetary masses, and $a/R_{\star}$ are taken from the Planetary Systems Composite Parameters table (PSCompPars [^3] https://exoplanetarchive.ipac.caltech.edu/cgi-bin/TblView/nph-tblView?app=ExoTbls&config=PSCompPars, 10.26133/NEA2) at NASA Exoplanet Archive (Christiansen2025); when a value is missing there, we adopt the entry from the Encyclopaedia of exoplanetary systems [^4] https://exoplanet.eu/catalog/ as a fallback.
@@
Among stellar-obliquity measurement techniques, the Rossiter-McLaughlin method is the most widely used and provides the majority of $\lambda$ measurements (WinnFabrycky2015; Triaud2018; Albrecht2022). Because different techniques introduce different selection effects (e.g., spot-crossing is more sensitive to aligned systems, while gravity-darkening is usually significant only for highly misaligned systems, Siegel2023; Albrecht2022), we retain only RM-based measurements, giving priority in the following order: classical RM, Doppler shadow (Albrecht2007; Collier2010; Zhou2016; Johnson2017), reloaded RM (Cegla2016ReloadedRM), and RM Revolution (Bourrier2021RMrevolutions). For all retained systems, we adopt the sky-projected obliquity $\lambda$ and $T_{\rm eff}$  from TEPCat [^5] https://www.astro.keele.ac.uk/jkt/tepcat/obliquity.html,
@@
Stellar multiplicity has long been theorized as a potential driver of spin-orbit misalignments (Holman1997; Wu2003; Fabrycky2007; Naoz2012; Batygin2012). Recent obliquity studies confirm that warm Jupiters tend to be aligned in single-star systems (Rice2022WJs_Aligned; WangX2024; Espinoza2023) yet are frequently (and sometimes famously, e.g., HD 80606; Winn2009; TIC 241249530; Gupta2024) misaligned in binaries; massive planets show a similar single-versus-binary contrast (Rusznak2025). Motivated by this, we identify and flag binaries or higher-order multiples in our obliquity sample using Schwarz2016 and Fontanive2021. Similar to subsection II.1, we examined the binarity of each system following the method of Badry2021, but extended it to include faint nearby stars ($G>18$). We also classified systems with RUWE >> 1.4 as binary candidates. Moreover, we only include systems with stellar ages >> 200 Myr and $\log g_{\star}$ >> 3.75 to exclude young and evolved stars, which is consistent with the construction of the $v\sin{i_{\star}}$  sample.
@@
Figure 3: The $\lambda$ distributions for hot-Jupiter ($3\times 10^{-4}\leq m_{\rm p}/M_{\star}\leq 2\times 10^{-3}$, $a/R_{\star}$<<10), sub-Saturn ($3\times 10^{-5}\leq m_{\rm p}/M_{\star}\leq 3\times 10^{-3}$), warm-Jupiter ($3\times 10^{-4}\leq m_{\rm p}/M_{\star}\leq 2\times 10^{-3}$, $a/R_{\star}$>>11), compact multiplanet (period ratio of adjacent planets, $P_{i+1}/P_{i}<6$), massive-planet systems ($2\times 10^{-3}\leq m_{\rm p}/M_{\star}$), and eccentric Jupiters ($3\times 10^{-4}\leq m_{\rm p}/M_{\star}\leq 2\times 10^{-3}$, $e>0.1$), along with the Kraft break and its associated uncertainty derived in this work. In terms of $T_{\rm eff}$  coverage, only the hot-Jupiter systems have sufficient measurements (>10>10) in the hot-star regime. Sub-Saturns with nearby companions are marked in blue. The predicted relation between $T_{\rm eff}$  and $\lambda$ from secular resonance crossing (Petrovich2020; Dugan2025) is shown as the light-blue shaded region.
III $T_{\rm eff}$  Boundary Determination